OpenStudy (anonymous):

Find all points (i.e. all x-values) between 0 and 2pi where the line tangent to the graph of y =sinx/2+cosx is horizontal

OpenStudy (tkhunny):

Have you considered the 1st Derivative?

OpenStudy (anonymous):

1st derivative would be -(cot x) (cosx/-sinx) how would I get the points from that?

OpenStudy (tkhunny):

If that is the correct 1st derivative, which I didn't check, you determine where it is zero.

OpenStudy (anonymous):

well the derivative of sinx is cosx and the derivative of any constant is always 0 and the derivative of cosx is -sinx and we know that cotx =cosx/sinx

OpenStudy (anonymous):

I am not sure how I would determine where it would be zero..

OpenStudy (anonymous):

horizontal tg = 0(no variation)

OpenStudy (tkhunny):

No. That will not do. You need the Quotient Rule. I get the 1st Derivative as \(\dfrac{2\cos(x)+1}{(2+\cos(x))^{2}}\). After that, simply concern yourself with the numerator being zero, unless the denominator happens to be zero in the same place.

OpenStudy (tkhunny):

This assumes your original expression was \(\dfrac{\sin(x)}{2+\cos(x)}\). Was it?

OpenStudy (anonymous):



OpenStudy (anonymous):

yes tkhunny that was the original equation

OpenStudy (tkhunny):

Then I gave the correct 1st Derivative and you should use parentheses to clarify your meaning. Remember your Order of Operations.

OpenStudy (tkhunny):

gtg sorry.

OpenStudy (anonymous):

rapahael can you still help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so do I need to set the numerator =0

OpenStudy (anonymous):

yes, because if the tangent line is horizontal, his angular coefficient is zero

OpenStudy (anonymous):

so then I end up with cosx=-1/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so is that my only x value , -1? ( my reasoning behind that is that cos of 60degrees=1/2) so then y x value is -1?

OpenStudy (anonymous):



OpenStudy (anonymous):

pi-pi/3,pi+pi/3

OpenStudy (anonymous):

@trusev1 get it?

OpenStudy (anonymous):

I get that but it is asking for the x-values that are at the line tangent is horizontal. so we got the derivative and even from your picture both those points would have the same x-value.

OpenStudy (anonymous):

no. my figure isnt the graph of your question,it's just the trigonometric circle